# Beyond Single-Event Extraction: Towards Efficient Document-Level Multi-Event Argument Extraction

### v1

@@ Abstract
Recent mainstream event argument extraction methods process each event in isolation, resulting in inefficient inference and ignoring the correlations among multiple events. To address these limitations, here we propose a multiple-event argument extraction model DEEIA (Dependency-guided Encoding and Event-specific Information Aggregation), capable of extracting arguments from all events within a document simultaneously. The proposed DEEIA model employs a multi-event prompt mechanism, comprising DE and EIA modules. The DE module is designed to improve the correlation between prompts and their corresponding event contexts, whereas the EIA module provides event-specific information to improve contextual understanding. Extensive experiments show that our method achieves new state-of-the-art performance on four public datasets (RAMS, WikiEvents, MLEE, and ACE05), while significantly saving the inference time compared to the baselines. Further analyses demonstrate the effectiveness of the proposed modules.
@@ 1 Introduction
Document-level event argument extraction (EAE) is a key process within Information Extraction Hobbs (2010); Grishman (2015), focused on identifying event-related arguments and their respective roles in document-level texts. Recently, the leading-edge methods for this task delve into prompt-based techniques Ma et al. (2022); Hsu et al. (2023), due to their great generalizability and competitive performance. Figure 1 (a) presents an example demonstrating EAE utilizing a prompt-based approach. Regarding the event $e_{0}$ triggered by “bombarding”, the approach defines a corresponding event prompt and identifies arguments: “government” as the killer, “a number of areas” as the victim, and “shelling” as the instrument.
@@
Mainstream EAE works Zhou et al. (2024); Liu et al. (2023c); Ren et al. (2023); Liu et al. (2023a) can only process one event at a time. When encountering documents containing multiple events, the limitations of these single-event argument extraction (Single-EAE) methods become evident. (1) As shown in Figure 1 (b), for a document containing multiple events, Single-EAE methods have to perform numerous iterations to extract event arguments for all events, which process the same document text repeatedly, leading to inefficient extraction. (2) Single-EAE methods fail to capture the beneficial event correlations among multiple events He et al. (2023); Zeng et al. (2022). Figure 1 (b) illustrates the argument overlapping phenomenon which reflects the semantic correlations among events. The event Life.death and event Life.surrender share the argument “government” and there exists a strong event correlation between these two events. However, Single-EAE methods cannot utilize such correlations.
@@ 2 Related Work
Recently, there has been an increasing interest in the task of document-level Event Argument Extraction Wang et al. (2022); Liu et al. (2023b); Yang et al. (2023), a crucial component within the domain of Event Extraction  Ren et al. (2022); Yang et al. (2021). Current methods for document-level EAE can be classified into four main categories: (1) Span-based methods, which identify candidate spans and subsequently predict their roles Zhang et al. (2020); Yang et al. (2023); Liu et al. (2017); Zhang et al. (2020); Liu et al. (2023c). (2) Generation-based methods Li et al. (2021); Du et al. (2021); Wei et al. (2021); Huang et al. (2023) utilizing generative PLMs, such as BART Lewis et al. (2019), to sequentially produce all arguments for the designated event. (3) Prompt-based methods Ma et al. (2022); He et al. (2023); Nguyen (2023), which use slotted prompts and leverage a generative slot-filling approach for argument extraction. (4) Large language model methods. Recently, some work Zhang et al. (2024); Zhou et al. (2023) has attempted to explore to utilize large language models for EAE tasks, but the performance falls short of expectations. And the time and cost of inference are relatively high. Among them, prompt-based methods have been demonstrated superior generalizability and competitive performance Hsu et al. (2023).
However, these EAE methods are Single-EAE methods, which can only process one event at a time. Recently, prompt-based EAE method, TabEAE He et al. (2023) aims to capture event co-occurrence Zeng et al. (2022) and trains the model on multi-event scheme. However, TabEAE requires separately processing prompts for different events, which is highly time-consuming. In this paper, we propose to extract event arguments concurrently, which significantly improves the efficiency of the EAE task in multi-event documents.
@@ 3.4 Span Selection
DEEIA(Ours) RoBERTa-l 58.0 53.4 71.8 67.0 75.4 74.5
@@ B.3 Experimental Results on ACE05
Table 6: Comparison of performance on ACE05 test set. * means we rerun their code based on their experimental settings. Bold and underline indicate the best and second-best experimental results.
@@ C.5 Comparison with Large Language Models
Large language models (LLMs) have garnered substantial interest and attention from researchers, highlighting their extensive applicability across a wide array of Information Extraction (IE) tasks Xu et al. (2023); Li et al. (2023b); Zhang et al. (2024). In this paper, we make a comparison with the recent state-of-the-art LLM-based approach presented in the work Zhou et al. (2023), which utilizes LLMs for the EAE task. We report their experimental results in Table 8. The experiments are conducted on three prominent large language models: text-davinci-003 Ouyang et al. (2022), gpt-3.5-turbo Ouyang et al. (2022) and GPT-4 Achiam et al. (2023). These models are accessed via the public APIs from OpenAI’s services [^7] https://openai.com/api/. As shown in Table 8, compared to the supervised learning models, LLMs still show a significant performance gap in the EAE task. Additionally, the operational costs of large models are inherently high. Our approach outperforms LLMs in terms of efficiency, cost, and effectiveness in the document-level EAE task.

### v2

@@ Abstract
Recent mainstream event argument extraction methods process each event in isolation, resulting in inefficient inference and ignoring the correlations among multiple events. To address these limitations, here we propose a multiple-event argument extraction model DEEIA (Dependency-guided Encoding and Event-specific Information Aggregation), capable of extracting arguments from all events within a document simultaneously. The proposed DEEIA model employs a multi-event prompt mechanism, comprising DE and EIA modules. The DE module is designed to improve the correlation between prompts and their corresponding event contexts, whereas the EIA module provides event-specific information to improve contextual understanding. Extensive experiments show that our method achieves new state-of-the-art performance on four public datasets (RAMS, WikiEvents, MLEE, and ACE05), while significantly saving the inference time compared to the baselines. Further analyses demonstrate the effectiveness of the proposed modules. Our implementation is available at https://github.com/LWL-cpu/DEEIA.
@@ 1 Introduction
Document-level event argument extraction (EAE) is a key process within Information Extraction Hobbs (2010); Grishman (2015); Xia et al. (2022), focused on identifying event-related arguments and their respective roles in document-level texts. Recently, the leading-edge methods for this task delve into prompt-based techniques Ma et al. (2022); Hsu et al. (2023), due to their great generalizability and competitive performance. Figure 1 (a) presents an example demonstrating EAE utilizing a prompt-based approach. Regarding the event $e_{0}$ triggered by “bombarding”, the approach defines a corresponding event prompt and identifies arguments: “government” as the killer, “a number of areas” as the victim, and “shelling” as the instrument.
@@
Mainstream EAE works Zhou et al. (2024); Liu et al. (2023c); Ren et al. (2023); Liu et al. (2023a) can only process one event at a time. When encountering documents containing multiple events, the limitations of these single-event argument extraction (Single-EAE) methods become evident. (1) As shown in Figure 1 (b), for a document containing multiple events, Single-EAE methods have to perform numerous iterations to extract event arguments for all events, which process the same document text repeatedly, leading to inefficient extraction. (2) Single-EAE methods fail to capture the beneficial event correlations among multiple events He et al. (2023); Zeng et al. (2022b). Figure 1 (b) illustrates the argument overlapping phenomenon which reflects the semantic correlations among events. The event Life.death and event Life.surrender share the argument “government” and there exists a strong event correlation between these two events. However, Single-EAE methods cannot utilize such correlations.
@@ 2 Related Work
Recently, there has been an increasing interest in the task of document-level Event Argument Extraction Wang et al. (2022); Liu et al. (2023b); Yang et al. (2023), a crucial component within the domain of Event Extraction  Ren et al. (2022); Yang et al. (2021). Current methods for document-level EAE can be classified into four main categories: (1) Span-based methods, which identify candidate spans and subsequently predict their roles Zhang et al. (2020); Yang et al. (2023); Liu et al. (2017); Zhang et al. (2020); Liu et al. (2023c). (2) Generation-based methods Li et al. (2021); Du et al. (2021); Wei et al. (2021); Huang et al. (2023) utilizing generative PLMs, such as BART Lewis et al. (2019), to sequentially produce all arguments for the designated event. (3) Prompt-based methods Ma et al. (2022); He et al. (2023); Nguyen (2023); Wang et al. (2024), which use slotted prompts and leverage a generative slot-filling approach for argument extraction. (4) Large language model methods. Recently, some work Zhang et al. (2024b); Zhou et al. (2023a) has attempted to explore to utilize large language models for EAE tasks, but the performance falls short of expectations. And the time and cost of inference are relatively high. Among them, prompt-based methods have been demonstrated superior generalizability and competitive performance Hsu et al. (2023).
However, these EAE methods are Single-EAE methods, which can only process one event at a time. Recently, prompt-based EAE method, TabEAE He et al. (2023) aims to capture event co-occurrence Zeng et al. (2022b) and trains the model on multi-event scheme. However, TabEAE requires separately processing prompts for different events, which is highly time-consuming. In this paper, we propose to extract event arguments concurrently, which significantly improves the efficiency of the EAE task in multi-event documents.
@@ 3.4 Span Selection
DEEIA(Ours) RoBERTa-l 58.0 53.4 71.8 67.0 75.2 74.3
@@ 7 Limitations
Acknowledgements
This work was supported by the National Science Foundation of China under Grant 62106038, and in part by the Sichuan Science and Technology Program under Grant 2023YFG0259.
@@ B.3 Experimental Results on ACE05
Table 6: Comparison of performance on ACE05 test set. * means we rerun their code based on their experimental settings. Bold indicates the best experimental results.
@@ C.5 Comparison with Large Language Models
Large language models (LLMs) have garnered substantial interest and attention from researchers, highlighting their extensive applicability across a wide array of tasks, such as text classification (Chen et al., 2021; Luo et al., 2024), dialogue (Zhang et al., 2023, 2024a), offensive language detection (Zhou et al., 2023c, b), graph tasks Zeng et al. (2023, 2022a), and in particular, the formation Extraction (IE) tasks Xu et al. (2023); Li et al. (2023b); Zhang et al. (2024b); Zhou et al. ; Liu et al. (2022, 2024); Xia et al. (2024a, b). In this paper, we make a comparison with the recent state-of-the-art LLM-based approach presented in the work Zhou et al. (2023a), which utilizes LLMs for the EAE task. We report their experimental results in Table 8. The experiments are conducted on three prominent large language models: text-davinci-003 Ouyang et al. (2022), gpt-3.5-turbo Ouyang et al. (2022) and GPT-4 Achiam et al. (2023). These models are accessed via the public APIs from OpenAI’s services [^7] https://openai.com/api/. As shown in Table 8, compared to the supervised learning models, LLMs still show a significant performance gap in the EAE task. Additionally, the operational costs of large models are inherently high. Our approach outperforms LLMs in terms of efficiency, cost, and effectiveness in the document-level EAE task.